# Locality Alignment Improves Vision-Language Models

### v1

@@ Abstract
Vision language models (VLMs) have seen growing adoption in recent years, but many still struggle with basic spatial reasoning errors. We hypothesize that this is due to VLMs adopting pre-trained vision backbones, specifically vision transformers (ViTs) trained with image-level supervision and minimal inductive biases. Such models may fail to encode the class contents at each position in the image, and our goal is to resolve this by ensuring that the vision backbone effectively captures both local and global image semantics. Our main insight is that we do not require new supervision to learn this capability – pre-trained models contain significant knowledge of local semantics that we can extract and use for scalable self-supervision. We propose a new efficient post-training stage for ViTs called locality alignment and a novel fine-tuning procedure called MaskEmbed that uses a masked reconstruction loss to learn semantic contributions for each image patch. We first evaluate locality alignment with a vision-only benchmark, finding that it improves a model’s performance at a patch-level semantic segmentation task, especially for strong backbones trained with image-caption pairs (e.g., CLIP and SigLIP). We then train a series of VLMs with and without locality alignment, and show that locality-aligned backbones improve performance across a range of benchmarks, particularly ones that involve spatial understanding (e.g., RefCOCO, OCID-Ref, TallyQA, VSR, AI2D). Overall, we demonstrate that we can efficiently learn local semantic extraction via a locality alignment stage, and that this procedure complements existing VLM training recipes that use off-the-shelf vision backbones.
@@ 1 Introduction
Auto-regressive VLMs are an exciting new type of model that emerged in the last couple years and has seen growing adoption (Alayrac et al., 2022). They are more flexible than previous multi-modal image-text models (Karpathy and Fei-Fei, 2015, Radford et al., 2021), leverage the reasoning abilities and open-ended nature of pre-trained language models (LMs) (Touvron et al., 2023, Jiang et al., 2023, Zheng et al., 2023), and have the potential to subsume many visual tasks that can be expressed in natural language and interwoven images (Lu et al., 2022, Chen et al., 2022a, Gupta et al., 2022).
However, current VLMs make a range of basic perceptual errors and struggle in particular with spatial understanding. Multiple recent works document such failures (Tong et al., 2024b, Rahmanzadehgervi et al., 2024), and weaknesses can be seen through benchmarks focused on object localization (Kazemzadeh et al., 2014, Wang et al., 2021), counting (Acharya et al., 2019) and relational question-answering (Liu et al., 2023a). Data limitations are part of the problem, because LMs might not fully exploit visual features without sufficient joint training. But we suspect that another issue is how these models leverage pre-trained vision backbones: the most popular current ViTs are trained with image-level supervision and minimal spatial inductive biases (e.g., CLIP and SigLIP; Radford et al. 2021, Zhai et al. 2023b), so they may fail to encode the necessary information for spatial reasoning. Ideally, we want a ViT whose representation is sufficient to predict class contents not only for the entire image but also for each region, which we refer to as encoding local semantics. Since most VLM training recipes either freeze or only partially train the ViT backbone (Liu et al., 2023c, Karamcheti et al., 2024, Laurençon et al., 2024, Lu et al., 2024, Bai et al., 2023), and because it may be difficult to learn local semantics during joint fine-tuning without extensive multi-modal data, we reason that it would help to use a ViT that better captures these rich image details.
Our goal in this work is to train a vision backbone that matches the best existing models in global image understanding (Radford et al., 2021, Zhai et al., 2023b) but that also encodes local semantics. We reason that disentangling where semantics arise in an image provides necessary information for certain downstream tasks, and sacrifices nothing if local semantics can collectively provide rich global image understanding. However, learning such a backbone is challenging due to limitations in current training approaches: for example, scalable objectives like CLIP offer only image-level supervision (Radford et al., 2021), semantic segmentation datasets contain relatively few images (Lin et al., 2014, Zhou et al., 2019, Gupta et al., 2019), and densely self-supervised methods like MAE and BEiT lack rich semantics (He et al., 2022, Bao et al., 2021).
@@
Since we bypass the need to train from scratch, we view this as a post-training stage for ViTs that we call locality alignment (Figure 1). The goal of this training stage is to take the set of concepts that an existing model is trained to recognize, and localize them by disentangling where they occur in an image. Our approach can be applied to any strong model trained with image-level supervision (e.g., CLIP, SigLIP, MoCo), leverages self-supervision instead of requiring costly human annotations, and has relatively low computational cost compared to pre-training. Our experiments focus on improving the performance of VLMs, but locality alignment may also prove useful for other applications.
To verify the effectiveness of locality alignment, we conduct both a vision-centric evaluation and a vision-language evaluation where we compare VLMs trained with different vision backbones. In our first set of experiments, we want to test whether locality-aligned ViTs encode what’s where in an image, and we measure this via a simple probing benchmark: we cast existing semantic segmentation datasets as a patch-wise multi-label classification problem (e.g., MSCOCO; Lin et al. 2014), and we find that locality alignment improves the performance of various backbones trained with image-level supervision, particularly language-supervised models like CLIP and SigLIP (Radford et al., 2021, Zhai et al., 2023b). Next, our main set of vision-language experiments compare a series of VLMs trained with and without locality alignment. We train our models using the recently released Prismatic library (Karamcheti et al., 2024) and with the strongest current ViT backbones, and we find that locality alignment improves performance across a range of benchmarks, particularly those that involve spatial reasoning (e.g., RefCOCO, OCID-Ref, TallyQA, VSR, AI2D). Through these experiments, we find that the best models for VLMs are reliably improved by locality alignment.
@@ 2 Related Work
ViT pre-training. There are many ways to pre-train ViTs, including strongly supervised approaches like image classification (Dosovitskiy et al., 2020), language-supervised objectives like CLIP and SigLIP (Radford et al., 2021, Yu et al., 2022, Zhai et al., 2023b, Tschannen et al., 2023), and various self-supervised tasks like BERT-style masked image modeling (Bao et al., 2021, He et al., 2022), augmentation-invariance (Chen et al., 2020b, Caron et al., 2021) and auto-regressive pixel generation (Chen et al., 2020a, El-Nouby et al., 2024). Pre-trained vision models are often adapted to downstream tasks, including semantic segmentation, object detection and depth estimation (Li et al., 2022b, Birkl et al., 2023, Kirillov et al., 2023), but training data for these tasks is typically scarce. Among these various training approaches, language-supervised models have proved most effective for VLMs in recent studies (Karamcheti et al., 2024, McKinzie et al., 2024, Tong et al., 2024a). Our work is motivated by a lack of training objectives with large-scale, dense and semantically rich supervision. We review existing pre-training approaches in more detail in Appendix A.
@@
ViT local feature extraction. Several works have noted CLIP’s lack of localized features in the context of downstream dense prediction tasks (Zhong et al., 2022, Rao et al., 2022, Xu et al., 2022, Wu et al., 2024). Other works have shown that ViTs learn to associate nearby patches (Dosovitskiy et al., 2020, Raghu et al., 2021, Jelassi et al., 2022), but this is distinct from encoding local semantics in their outputs. Some have proposed hybrid ViTs that reintroduce inductive biases from CNNs (Liu et al., 2021, Wu et al., 2021, d’Ascoli et al., 2021), but we improve the original ViT’s local feature extraction without sacrificing expressive power. The works most closely related to ours are RegionCLIP (Zhong et al., 2022), CLIPSelf (Wu et al., 2024) and LocCa (Wan et al., 2024). RegionCLIP fine-tunes CLIP with synthetically labeled region-text pairs, which avoids human annotation but suffers from noisy caption matching. CLIPSelf fine-tunes CLIP to reconstruct features of random image sub-crops, which is similar to our approach but specifically intended for zero-shot semantic segmentation; this difference in goals leads to suboptimal localization under probing, as we show in Section 4. LocCa is trained to auto-regressively predict synthetic image captions from OWL-ViT (Minderer et al., 2022), which is itself a CLIP model fine-tuned on dense object annotations. Compared to LocCa, our approach requires significantly less compute, does not require any extra human annotations, and can be flexibly applied to any pre-trained model. [^1] We are unable to compare to LocCa (Wan et al., 2024) due to a lack of public checkpoints.
@@
VLMs. We focus on the class of open-ended vision-augmented LMs, which includes early examples like Flamingo, OFA, BLIP and Llava (Alayrac et al., 2022, Wang et al., 2022, Li et al., 2022a, Liu et al., 2023c), and current frontier models like Claude 3.5 Sonnet, GPT-4o and Gemini 1.5 (OpenAI, , Anthropic, , Reid et al., 2024). The most common approach to building such models is to combine a pre-trained ViT and a pre-trained LM (Bai et al., 2023, Lu et al., 2024, Beyer et al., 2024), which leverages strong capabilities learned from each modality. Several recent works investigate how to best integrate visual features (Laurençon et al., 2024, McKinzie et al., 2024, Karamcheti et al., 2024, Tong et al., 2024a). Most use high-resolution variants of CLIP or SigLIP for their vision backbone (Radford et al., 2021, Zhai et al., 2023b) and either freeze or only partially train the ViT alongside the LM, which makes it important for the initial ViT to capture local semantics.
@@
VLM perceptual failures. VLMs are a diverse class of models with different interfaces and architectures, but many works have demonstrated perceptual errors across various types of multi-modal models (Thrush et al., 2022, Kamath et al., 2023, Yuksekgonul et al., 2023, Xu et al., 2024b). For the current generation of open-ended VLMs, perceptual flaws are apparent in benchmarks for counting, object localization, relational question-answering, object hallucination, and others like BlindTest (Rahmanzadehgervi et al., 2024) and MMMV (Tong et al., 2024b). Many of these tasks require spatial understanding, and we suspect that part of the problem is a failure to encode local image semantics. There are other ways to approach the issue, but an improved vision backbone composes with many of them: these include fusing features from multiple backbones (Karamcheti et al., 2024, Jain et al., 2024) or multiple image crops (Liu et al., 2024, Xu et al., 2024b), adding extra parameters for image processing (Tong et al., 2024a), and training with more data focused on spatial reasoning (Lu et al., 2022, Wang et al., 2023b, Peng et al., 2023, Xu et al., 2024a).
@@ 3.1 Masking is all you need
Consider a model trained to extract a rich global representation but no specific information for each image region, e.g., a CLIP image encoder (Radford et al., 2021). We want to use such a model to understand what’s where in the image, and we propose to do so with masking. A model that accurately represents global image contents will change its output in response to input masking, and we can exploit this to probe a model under different masked views and understand each patch’s contribution to the prediction. For example, comparing the output before and after masking a single patch provides information about that region’s contents (Zeiler and Fergus, 2014).
@@
This approach is a reasonable starting point, and it illustrates that pre-trained models contain latent knowledge of local semantics that can be extracted via masking. It also has a precedent in the literature: querying pre-trained models with masked images was one of the earliest approaches to zero-shot semantic segmentation (Xu et al., 2022), and this learning approach is the basis of certain interpretability methods (Jethani et al., 2021, Covert et al., 2022). However, we find that the additive approximation is limiting and not very effective in our experiments; this is because 1) patch semantics aren’t truly additive and the poor approximation causes us to lose information about each patch, 2) vector embeddings only allow us to reconstruct vector targets (e.g., the [CLS] token), which contain a small part of the pre-trained model’s information about the image. Our main approach described in the next section therefore generalizes this idea to learn richer patch embeddings.
@@ 3.2 Proposed approach
Compared to the earlier additive reconstruction approach (Section 3.1), MaskEmbed’s use of an expressive decoder helps compress more information into each patch embedding. This also differentiates our approach from CLIPSelf (Wu et al., 2024), which adopts a related objective but aggregates CLIP’s features by average-pooling within crop windows. We show the importance of this design choice in Section 4, where we also perform an ablation study to determine several hyperparameters for MaskEmbed. We remark that the main disadvantage of our approach is that our patch embeddings are less interpretable, because they lie in a different embedding space than the pre-trained model’s outputs; however, we reason that this is acceptable because our eventual use case is training a VLM that can learn how the entire representation encodes semantics.
@@ 3.3 Training data
MaskEmbed is supervised by the pre-trained model’s masked outputs, which means we can use any image dataset regardless of its annotations or lack thereof. Diverse data covering the pre-training distribution will help localize the broadest possible semantics, ideally including objects, backgrounds, textures, facial features, etc. We use ImageNet-1k and ImageNet-21k (hereafter IN1k and IN21k) (Deng et al., 2009) for all our experiments, which are relatively diverse and contain 1.2M and 12.6M images in our training sets. A promising idea that we leave to future work is using larger web-scraped image datasets like those used for contrastive learning (Schuhmann et al., 2022, Xu et al., 2023, Gadre et al., 2023, Fang et al., 2023a), which are even more diverse and could help learn strong localized text features that are less prominent in ImageNet.
Related to training data, we note that our approach only works as intended if the pre-trained model makes meaningful predictions with masked inputs. This can be ensured by pre-training with randomly dropped patches, which is performed for some but not all of the models in our experiments (He et al., 2022, Bao et al., 2021, Peng et al., 2022, Fang et al., 2024). Training or fine-tuning with random masking is often suggested in the interpretability literature (Frye et al., 2020, Covert et al., 2021, Jain et al., 2022) because masked images are out-of-distribution if the model was not trained with masking, but we do not explore this direction and instead rely on the fact that ViTs empirically behave reasonably under masking (Naseer et al., 2021).

### v2

@@ Abstract
Vision language models (VLMs) have seen growing adoption in recent years, but many still struggle with basic spatial reasoning errors. We hypothesize that this is due to VLMs adopting pre-trained vision backbones, specifically vision transformers (ViTs) trained with image-level supervision and minimal inductive biases. Such models may fail to encode the class contents at each position in the image, and our goal is to resolve this with a vision backbone that effectively captures both local and global image semantics. Our main insight is that we do not require new supervision to learn this capability – pre-trained models contain significant knowledge of local semantics that we can extract and use for scalable self-supervision. We propose a new efficient post-training stage for ViTs called locality alignment and a novel fine-tuning procedure called MaskEmbed that uses a masked reconstruction loss to learn semantic contributions for each image patch. We first evaluate locality alignment with a vision-only benchmark, finding that it improves a model’s performance at patch-level semantic segmentation, especially for strong backbones trained with image-caption pairs (e.g., CLIP and SigLIP). We then train a series of VLMs with and without locality alignment, and show that locality-aligned backbones improve performance across a range of benchmarks, particularly ones that involve spatial understanding (e.g., RefCOCO, OCID-Ref, TallyQA, VSR, AI2D). Overall, we demonstrate that we can efficiently learn local semantic extraction via a locality alignment stage, and that this procedure benefits VLM training recipes that use off-the-shelf vision backbones.
@@ 1 Introduction
Auto-regressive VLMs are an exciting new type of model that emerged in the last couple years and has seen growing adoption (Alayrac et al., 2022). They are more flexible than previous multi-modal image-text models (Karpathy & Fei-Fei, 2015; Radford et al., 2021), leverage the reasoning abilities and open-ended nature of pre-trained language models (LMs) (Touvron et al., 2023; Jiang et al., 2023; Zheng et al., 2023), and have the potential to subsume many visual tasks that can be expressed in natural language with interwoven images (Lu et al., 2022; Chen et al., 2022a; Gupta et al., 2022).
However, current VLMs make a range of basic perceptual errors and struggle in particular with spatial understanding. Multiple recent works document such failures (Tong et al., 2024b; Rahmanzadehgervi et al., 2024), and weaknesses can be seen through benchmarks focused on object localization (Kazemzadeh et al., 2014; Wang et al., 2021), counting (Acharya et al., 2019) and relational question-answering (Liu et al., 2023a). Data limitations are part of the problem, because LMs might not fully exploit visual features without sufficient joint training. But we suspect that another issue is how these models leverage pre-trained vision backbones: the most popular current ViTs are trained with image-level supervision and minimal spatial inductive biases (e.g., CLIP and SigLIP; Radford et al. 2021; Zhai et al. 2023b), so they may fail to encode the necessary information for spatial reasoning. Ideally, we want a ViT whose representation is sufficient to predict class contents not only for the entire image but also for each region, which we refer to as encoding local semantics. Since most VLM training recipes either freeze or only partially train the ViT backbone (Liu et al., 2023c; Karamcheti et al., 2024; Laurençon et al., 2024; Lu et al., 2024; Bai et al., 2023), and because it may be difficult to learn local semantics during joint fine-tuning without extensive multi-modal data, we reason that it would help to use a ViT that better captures these rich image details.
Our goal in this work is to train a vision backbone that matches the best existing models in global image understanding (Radford et al., 2021; Zhai et al., 2023b) but that also encodes local semantics. We reason that disentangling where semantics arise in an image provides necessary information for certain downstream tasks, and sacrifices nothing if local semantics collectively provide rich global image understanding. However, learning such a backbone is challenging due to limitations in current training approaches: for example, scalable objectives like CLIP offer only image-level supervision (Radford et al., 2021), semantic segmentation datasets contain relatively few images (Lin et al., 2014; Zhou et al., 2019; Gupta et al., 2019), and densely self-supervised methods like MAE and BEiT lack rich semantics (He et al., 2022; Bao et al., 2021).
@@
Since we do not require training from scratch, we view this as a post-training stage for ViTs that we call locality alignment (Figure 1). The goal of this training stage is to take the set of concepts that an existing model is trained to recognize, and localize them by disentangling where they occur in an image. Our approach can be applied to any strong model trained with image-level supervision (e.g., CLIP, SigLIP, MoCo), leverages self-supervision instead of requiring costly human annotations, and has relatively low computational cost compared to pre-training. Our experiments focus on improving the performance of VLMs, but locality alignment may also prove useful for other downstream tasks.
To verify the effectiveness of locality alignment, we conduct both a vision-centric evaluation and a vision-language evaluation where we compare VLMs trained with different vision backbones. In our first set of experiments, we want to test whether locality-aligned ViTs encode what’s where in an image, and we measure this via a simple probing benchmark: we cast existing semantic segmentation datasets as a patch-wise multi-label classification problem (e.g., MSCOCO; Lin et al. 2014) and find that locality alignment improves the performance of various backbones trained with image-level supervision, particularly language-supervised models like CLIP and SigLIP (Radford et al., 2021; Zhai et al., 2023b). Next, our main set of vision-language experiments compare a series of VLMs trained with and without locality alignment. We train our models using the recently released Prismatic library (Karamcheti et al., 2024) and with the strongest current ViT backbones, and we find that locality alignment improves performance across a range of benchmarks, particularly those that involve spatial reasoning (e.g., RefCOCO, OCID-Ref, TallyQA, VSR, AI2D). Through these experiments, we find that the best models for VLMs are reliably improved by locality alignment.
@@ 2 Related Work
ViT pre-training. There are many ways to pre-train ViTs, including strongly supervised approaches like image classification (Dosovitskiy et al., 2020), language-supervised objectives like CLIP and SigLIP (Radford et al., 2021; Yu et al., 2022; Zhai et al., 2023b; Tschannen et al., 2023), and various self-supervised tasks like BERT-style masked image modeling (Bao et al., 2021; He et al., 2022), augmentation-invariance (Chen et al., 2020b; Caron et al., 2021) and auto-regressive pixel generation (Chen et al., 2020a; El-Nouby et al., 2024). Pre-trained vision models are often adapted to downstream tasks, including semantic segmentation, object detection and depth estimation (Li et al., 2022b; Birkl et al., 2023; Kirillov et al., 2023), but training data for these tasks is typically scarce. Among these various training approaches, language-supervised models have proved most effective for VLMs in recent studies (Karamcheti et al., 2024; McKinzie et al., 2024; Tong et al., 2024a). Our work is motivated by a lack of training objectives with large-scale, dense and semantically rich supervision. We review existing pre-training approaches in more detail in Appendix A.
@@
ViT local feature extraction. Several works have noted CLIP’s lack of localized features in the context of downstream dense prediction tasks (Zhong et al., 2022; Rao et al., 2022; Xu et al., 2022; Wu et al., 2024). Other works have shown that ViTs learn to associate nearby patches (Dosovitskiy et al., 2020; Raghu et al., 2021; Jelassi et al., 2022), but this is distinct from encoding local semantics in their outputs. Some have proposed hybrid ViTs that reintroduce inductive biases from CNNs (Liu et al., 2021; Wu et al., 2021; d’Ascoli et al., 2021), but we improve the original ViT’s local feature extraction without sacrificing expressive power. The works most closely related to ours are RegionCLIP (Zhong et al., 2022), CLIPSelf (Wu et al., 2024) and LocCa (Wan et al., 2024). RegionCLIP fine-tunes CLIP with synthetically labeled region-text pairs, which avoids human annotation but suffers from noisy caption matching. CLIPSelf fine-tunes CLIP to reconstruct features of random image sub-crops, which is similar to our approach but specifically intended for zero-shot semantic segmentation; this difference in goals leads to suboptimal localization under probing, as we show in Section 4. LocCa is trained to auto-regressively predict synthetic image captions from OWL-ViT (Minderer et al., 2022), which is itself a CLIP model fine-tuned on dense object annotations. Compared to LocCa, our approach requires significantly less compute, does not require any extra human annotations, and can be flexibly applied to any pre-trained model. [^1] We are unable to compare to LocCa (Wan et al., 2024) due to a lack of public checkpoints.
@@
VLMs. We focus on the class of auto-regressive vision-augmented LMs, which includes early examples like Flamingo, OFA, BLIP and Llava (Alayrac et al., 2022; Wang et al., 2022; Li et al., 2022a; Liu et al., 2023c), and current frontier models like Claude 3.5 Sonnet, GPT-4o and Gemini 1.5 (OpenAI, ; Anthropic, ; Reid et al., 2024). The most common approach to building such models is to combine a pre-trained ViT and a pre-trained LM (Bai et al., 2023; Lu et al., 2024; Beyer et al., 2024), which leverages strong capabilities learned from each modality. Several recent works investigate how to best integrate visual features (Laurençon et al., 2024; McKinzie et al., 2024; Karamcheti et al., 2024; Tong et al., 2024a). Most use high-resolution variants of CLIP or SigLIP for their vision backbone (Radford et al., 2021; Zhai et al., 2023b) and either freeze or only partially train the ViT alongside the LM, which makes it important for the initial ViT to capture local semantics.
@@
VLM perceptual failures. VLMs are a diverse class of models with different interfaces and architectures, but many works have demonstrated perceptual errors across various types of multi-modal models (Thrush et al., 2022; Kamath et al., 2023; Yuksekgonul et al., 2023; Xu et al., 2024b). For the current generation of auto-regressive VLMs, perceptual flaws are apparent in benchmarks for counting, object localization, relational question-answering, object hallucination, and others like BlindTest (Rahmanzadehgervi et al., 2024) and MMMV (Tong et al., 2024b). Many of these tasks require spatial understanding, and we suspect that part of the problem is a failure to encode local image semantics. There are other ways to approach the issue, but an improved vision backbone composes with many of them: these include fusing features from multiple backbones (Karamcheti et al., 2024; Jain et al., 2024) or multiple image crops (Liu et al., 2024; Xu et al., 2024b), adding extra parameters for image processing (Tong et al., 2024a), and training with more data focused on spatial reasoning (Lu et al., 2022; Wang et al., 2023b; Peng et al., 2023; Xu et al., 2024a).
@@ 3.1 Masking is all you need
Consider a model trained to extract a rich global representation but no specific information for each image region, e.g., a CLIP image encoder (Radford et al., 2021). We want to use such a model to understand what’s where in the image, and we propose to do so with masking. A model that accurately represents global image contents will change its output in response to input masking, and we can exploit this to probe a model under different masked views and understand each patch’s contribution to the prediction. For example, comparing the output before and after masking a single patch provides information about that region’s contents (Zeiler & Fergus, 2014).
@@
This approach is a reasonable starting point, and it illustrates that pre-trained models contain latent knowledge of local semantics that can be extracted via masking. It also has a precedent in the literature: querying pre-trained models with masked images was one of the earliest approaches to zero-shot semantic segmentation (Xu et al., 2022), and this learning approach is the basis of certain interpretability methods (Jethani et al., 2021; Covert et al., 2022). However, we find that the additive approximation is limiting and not very effective in our experiments; this is because 1) patch semantics aren’t truly additive and the poor approximation causes us to lose information about each patch, 2) vector embeddings only allow us to reconstruct vector targets (e.g., the [CLS] token), which contain a small part of the pre-trained model’s information about the image. Our main approach presented in the next section therefore generalizes this idea to learn richer patch embeddings.
@@ 3.2 Proposed approach
Compared to the simpler additive reconstruction approach (Section 3.1), MaskEmbed’s use of an expressive decoder helps compress more information into each patch embedding. This also differentiates our approach from CLIPSelf (Wu et al., 2024), which adopts a related objective but aggregates CLIP’s features by average-pooling within crop windows. We show the importance of this design choice in Section 4, where we also perform an ablation study to determine several hyperparameters for MaskEmbed. We remark that the main disadvantage of our approach is that our patch embeddings are less interpretable, because they lie in a different embedding space than the pre-trained model’s outputs; however, we reason that this is acceptable because our eventual use case is training a VLM that can learn how the entire representation encodes semantics.
@@ 3.3 Training data
MaskEmbed is supervised by the pre-trained model’s masked outputs, which means we can use any image dataset regardless of its annotations or lack thereof. Diverse data covering the pre-training distribution will help localize the broadest possible semantics, ideally including many types of objects, backgrounds, textures, facial features, etc. We use ImageNet-1k and ImageNet-21k (hereafter IN1k and IN21k) (Deng et al., 2009) for all our experiments, which are relatively diverse and contain 1.2M and 12.6M images in our training sets. A promising idea that we leave to future work is using larger web-scraped image datasets like those used for contrastive learning (Schuhmann et al., 2022; Xu et al., 2023; Gadre et al., 2023; Fang et al., 2023a), which are even more diverse and could help learn strong localized text features that are less prominent in ImageNet.
Related to training data, we note that our approach only works as intended if the pre-trained model makes meaningful predictions with masked inputs. This can be ensured by pre-training with randomly dropped patches, which is performed for some but not all of the models in our experiments (He et al., 2022; Bao et al., 2021; Peng et al., 2022; Fang et al., 2024). Training or fine-tuning with random masking is often suggested in the interpretability literature (Frye et al., 2020; Covert et al., 2021; Jain et al., 2022) because masked images are out-of-distribution if the model was not trained with masking, but we do not explore this direction and instead rely on the fact that ViTs empirically behave reasonably under masking (Naseer et al., 2021). [^2] In Section B.4, we conduct initial experiments that suggest further gains if locality alignment is preceded by fine-tuning with randomly masked patches.